### v3

# Video-RAG: Visually-aligned Retrieval-Augmented Long Video Comprehension

Abstract
Existing large video-language models (LVLMs) struggle to comprehend long videos correctly due to limited context. To address this problem, fine-tuning long-context LVLMs and employing GPT-based agents have emerged as promising solutions. However, fine-tuning LVLMs would require extensive high-quality data and substantial GPU resources, while GPT-based agents would rely on proprietary models (e.g., GPT-4o). In this paper, we propose Video Retrieval-Augmented Generation (Video-RAG), a training-free and cost-effective pipeline that employs visually-aligned auxiliary texts to help facilitate cross-modality alignment while providing additional information beyond the visual content. Specifically, we leverage open-source external tools to extract visually-aligned information from pure video data (e.g., audio, optical character, and object detection), and incorporate the extracted information into an existing LVLM as auxiliary texts, alongside video frames and queries, in a plug-and-play manner. Our Video-RAG offers several key advantages: (i) lightweight with low computing overhead due to single-turn retrieval; (ii) easy implementation and compatibility with any LVLM; and (iii) significant, consistent performance gains across long video understanding benchmarks, including Video-MME, MLVU, and LongVideoBench. Notably, our model demonstrates superior performance over proprietary models like Gemini-1.5-Pro and GPT-4o when utilized with a 72B model.  [^1] Our code is available at https://github.com/Leon1207/Video-RAG-master.
1 Introduction
Figure 1: Illustration of two common approaches for understanding long videos, alongside our Video-RAG. Video-RAG provides a resource-efficient, training-free pipeline that is easily compatible with any LVLM. By leveraging RAG, it retrieves auxiliary texts for input, leading to notable performance enhancement.
With the advancements in Large Language Models (LLMs), numerous studies have been conducted to enhance their ability to comprehend and process videos [10, 14, 16, 48, 22, 45, 17, 3, 2, 50, 21, 23, 15], collectively termed Large Video-Language Models (LVLMs). Although current LVLMs have demonstrated promising performance in understanding short videos, effective comprehension of extremely long videos continues to be a major challenge.
To address this challenge, recent studies [47, 43, 33, 39, 53] have sought to extend the reasoning context length of LVLMs, essentially finetuning long-context LVLMs for long video understanding. LongVA [47] first introduces increasing the token capacity of an LLM and transferring its long-context comprehension capabilities to video data. However, training such a model requires pre-training on an extended corpus, and often there are distribution shifts between deployment videos and finetuning videos. As demonstrated in Video-MME [6], LongVA declines when increasing the video frame sampling rate from 128 to 384 (52.6% → 51.8%). This outcome suggests that simply increasing the number of sampled frames not only leads to information redundancy but also imposes additional challenges for the model to handle complex reasoning. Retrieval-Augmented Generation [12] (RAG) is a technique that enhances generative tasks by retrieving relevant documents from an external corpus, thus improving response quality in LLMs. Recent studies have begun exploring the integration of RAG with video-based tasks [1, 25, 46], employing tools to process videos in long contexts and sending them to a proprietary model for generation, which is known as the GPT-based Agent method. However, they come with serval limitations. First, most of them process long video content as plain text, subsequently utilizing the RAG mechanisms to retrieve relevant documents for LLMs. Therefore, they lack alignment with the visual context of the video, resulting in a loss of critical visual information. Second, they are often resource-intensive in multi-turn interactions and typically require powerful LLMs to function as the driving force, thus limiting their flexibility and generative capabilities. Executing the whole Video-MME [6] using VideoAgent [4] requires approximately 20 days and incurs a substantial consumption of GPT-4o API tokens.
Figure 2: Comparison of the performance of Video-RAG with LLaVA-Video-72B [49], Gemini-1.5-Pro [31], and GPT-4o [28] across various benchmarks, including the sub-tasks from Video-MME [6] (here we focus only on those that outperform Gemini-1.5-Pro), LongVideoBench [41], and MLVU [52] benchmarks.
In this study, we propose Video-RAG, an effective RAG pipeline that can be seamlessly integrated with any LVLM. Specifically, instead of simply increasing the number of sampled video frames, we propose to replace the corresponding extended visual tokens with auxiliary texts extracted from pure video data by invoking open-source foundation models, such as optical character recognition (OCR), automatic speech recognition (ASR), and object detection. These auxiliary texts are more aligned with the visual context while providing additional information beyond the visual data, as demonstrated in [18, 4]. Besides dealing with the context windows limit of LVLMs, we employ RAG in Video-RAG to filtering auxiliary texts, ensuring their relevance to the user’s query in the text embedding space. As sampled visual context often lacks explicit alignment with the instructions, the auxiliary texts can facilitate cross-modality alignment while reducing the modality divide. As illustrated in Figure 6, with Video-RAG, the retrieved auxiliary texts help guide the LVLM to pay more attention to the query-relevant keyframes, while simultaneously facilitating cross-modality alignment between query and keyframes. In this framework, an LVLM serves as the central component of Video-RAG, processing visual tokens to preserve detailed visual context and minimize potential information loss. Moreover, the retrieval process is parallelly executed in a single operation, ensuring the efficiency of our pipeline.
We evaluate Video-RAG across several long video benchmarks, including Video-MME [6], MLVU [52], and LongVideoBench [41]. By applying the Video-RAG to six distinctive open-source LVLMs, we achieve an average performance improvement of 8.0% on Video-MME with only 2.0K text tokens addition (equal to 14 frames in most configuration) per case, while beating the proprietary LVLM Gemini-1.5-Pro [31] when integrated with the 72B model, as shown in Figure 7. Applying Video-RAG to a 7B LVLM only requires an additional 8GB of inference GPU memory and approximately 5 seconds of inference time per case.
In summary, our contributions are as follows:
  • We integrate RAG into open-source LVLMs: [^2] While some methods [25, 46, 1] use RAG system for video tasks, they convert the video data fully into text while also relying on proprietary, non-open-source models. These approaches may result in the loss of visual information and lead to significant resource and time consumption. Video-RAG incorporates three types of visually-aligned auxiliary texts (OCR, ASR, and object detection) processed by external tools and retrieved via RAG, enhancing the LVLM. It’s implemented using completely open-source tools, without the need for any commercial APIs.
  • We design a versatile plug-and-play RAG-based pipeline for any LVLM: Video-RAG offers a training-free solution for a wide range of LVLMs, delivering performance improvements with minimal additional resource requirements.
  • We achieve proprietary-level performance with open-source models: Applying Video-RAG to a 72B open-source model yields state-of-the-art performance in Video-MME, surpassing models such as Gemini-1.5-Pro.
2 Related Work
Figure 3: The framework of our Video-RAG pipeline that contains three key phases. In the query decouple phase, the LVLM is prompted to generate a retrieval request for auxiliary texts. Next, in the auxiliary text generation and retrieval phase, the video is processed in parallel to extract three types of textual information (OCR, ASR, and object detection), and the relevant text is retrieved as the auxiliary text. Finally, in the integration and generation phase, auxiliary texts are combined with the query and the video to generate the response.
2.1 Large Video-Language Models
With the rapid advancement of large language models (LLMs), there has been increasing interest in developing generalist video models capable of handling a wide range of video-related tasks. Video-ChatGPT [26] extracts features from individual frames and aggregates them through both spatial and temporal pooling operations. VideoChat [14] encodes videos by generating both textual descriptions and video appearance embeddings. Video-LLaVA [16] aligning image and video encoders during a pre-processing phase, using a shared projector to map the encoded representations into a common language space. LLaVA-NeXT-Video [48] extends LLaVA-NeXT [20] by fine-tuning the model specifically on video data. Despite their contributions, these approaches face challenges when processing detailed and long-length videos, primarily due to the limited number of frames sampled for analysis.
2.2 Long-context Large Video-Language Models
Recent approaches have sought to expand the context window size to enhance detailed video understanding. LongVA [47] and Long-LLaVA [43] address this by continuously training LLMs on extended textual data, to transfer their long-text comprehension capabilities to video processing. INTP [33] introduces a video token rearrangement technique while proposing a training-free method for extending the LLM context window, allowing LVLMs to process increased visual tokens. However, these methods face challenges in striking a balance between the high computational costs associated with sampling video frames and the limited performance improvements achieved. Due to the inherent redundancy in video content and constraints on model capacity, performance degradation may occur when the number of sampled frames surpasses a certain threshold.
2.3 GPT-based Agent Video Understanding
Initial efforts [27, 44, 40, 7, 36] have employed LLMs to interact with tools to process visual information as structured long context for question-answering. MM-VID [18] enhances long video understanding by aligning video frames with corresponding text descriptions. VLog [19] leverages multimodel pre-trained models to capture and interpret visual and audio information, summarizing it into documents for video comprehension. VideoAgent [4], DrVideo [25], and OmAgent [46] integrate multimodal inputs and enable dynamic querying of video segments to support long video reasoning tasks. However, these methods take an extremely long time to process videos while relying on proprietary models (e.g., GPT-4o), thus limiting their efficiency and adaptability to other open-source frameworks.
3 Method
We propose a novel, training-free pipeline for large video-language models (LVLMs), named Video-RAG, which can be integrated into any LVLM. As illustrated in Figure 3, our pipeline comprises three key phases: (i) Query Decouple: In this phase, the user’s query is decomposed into a retrieval request aimed at extracting auxiliary texts from the target video. (ii) Auxiliary Text Generation & Retrieval: Multiple auxiliary texts are generated from the queried video in parallel. Then, the retrieval request is used to obtain relevant external information. (iii) Integration and Generation: This phase integrates the retrieved auxiliary texts with the user’s query, feeding this combined input into the LVLMs to generate the final response.
3.1 Large Video-Language Model
Given a video $\bm{\mathrm{V}}$, a frame sampler first sample $N$ frames $\bm{\mathrm{F}}$. Most existing methods uniformly sample frames from a video for both effectiveness and simplicity. Then, video features are extracted as $\bm{\mathrm{F}_{v}}=\texttt{VisualEnc}(\bm{\mathrm{F}})$, where VisualEnc is an image-based visual encoder, such as CLIP-L [29]. Finally, the video features $\bm{\mathrm{F}_{v}}$ and the user’s query $\bm{\mathrm{Q}}$ are fed into the LVLM to generate an output $\bm{\mathrm{O}}$:
$\bm{\mathrm{O}}=\texttt{LVLM}(\bm{\mathrm{F}_{v}},\bm{\mathrm{Q}})$ (1)
3.2 Query Decouple
In this phase, upon receiving a user’s query about the video, the LVLM begins by decoupling the query and generating retrieval requests, denoted as $\bm{\mathrm{R}}$, for auxiliary texts. During this phase, the LVLM processes only textual information, without access to video frames, and the output requests are formatted in JSON. We prompt the LVLM using a decoupling prompt $\bm{\mathrm{P}}$ to generate the following retrieval requests as necessary: (i) $\bm{\mathrm{R}}_{asr}$: Requests about automatic speech recognition, to extract audio information from the video that may pertain to the query. (ii) $\bm{\mathrm{R}}_{det}$: Requests for identifying physical entities within the video that may assist in answering the query. (iii) $\bm{\mathrm{R}}_{type}$: Requests for details about the location, quantity, and relationships of the identified physical entities. These requests, which may be NULL (indicating that the corresponding information is not required), are then parsed and forwarded to the auxiliary text retrieval phase. The entire process can be formally described as:
$\bm{\mathrm{R}}=\texttt{LVLM}(\bm{\mathrm{P}},\bm{\mathrm{Q}}),\ \ \bm{\mathrm{R}}=\{\bm{\mathrm{R}}_{asr},\bm{\mathrm{R}}_{det},\bm{\mathrm{R}}_{type}\}$ (2)
3.3 Auxiliary Text Generation
In this phase, we first generate the auxiliary texts from the video and then retrieve them to assist the LVLMs according to the retrieval requests $\bm{\mathrm{R}}$. As the length of the video increases, the number of tokens generated from the processed data also grows, leading to an increase in redundant information. Additionally, current open-source models are constrained by the limited length of their context windows, which may prevent them from fully processing all auxiliary texts. To address this issue, we draw inspiration from Retrieval-Augmented Generation (RAG) [12], retrieving only the auxiliary texts relevant to the user’s query. Before retrieval, we construct the necessary databases from the given video in parallel. Specifically, we implement three distinct databases: the Optical Character Recognition (OCR) database, denoted as $DB_{ocr}$; the Automatic Speech Recognition (ASR) database, denoted as $DB_{asr}$; and the Object Detection (DET) database, denoted as $DB_{det}$.
OCR database. Current LVLM are still illusory in their ability to accurately recognize characters, and their performance often falls short compared to proprietary models. To better leverage the information contained in video frames and reduce hallucinations, we employ a proprietary OCR model to extract text from each sampled video frame. Specifically, we use EasyOCR [9] as our text recognition model and segmented the recognized texts on a per-frame basis, denoted as $\bm{\mathrm{T}}_{ocr}$. Subsequently, we implemented RAG by utilizing the advanced text encoding model Contriever [8] to encode the fetched OCR texts into text embeddings $\bm{\mathrm{E}}_{ocr}$. These embeddings are then stored in a database with the FAISS index [11], a library designed for efficient similarity search and clustering of dense vectors. The entire building process can be formally described as:
$\bm{\mathrm{T}}_{ocr}=\texttt{EasyOCR}(\bm{\mathrm{F}})$ (3)
$DB_{ocr}\xleftarrow{\ \texttt{FAISS}\ }\bm{\mathrm{E}}_{ocr}=\texttt{Contriever}(\bm{\mathrm{T}}_{ocr})$ (4)
ASR database. Audio information (e.g., subtitles) plays a crucial role in video comprehension, often providing additional context that may not be available through visual cues alone. To incorporate them, we first extract the raw audio $\bm{\mathrm{U}}$ from the video and then transcribe them into texts $\bm{\mathrm{T}}_{asr}$. Specifically, we use Whisper [30] as our audio transcription model. Since the recognized texts can be quite extensive, we chunk and encode them into a vector database, following the same procedure used to construct the OCR database. The building process can be formally described as:
$\bm{\mathrm{T}}_{asr}=\texttt{Whisper}(\bm{\mathrm{U}})$ (5)
$DB_{asr}\xleftarrow{\ \texttt{FAISS}\ }\bm{\mathrm{E}}_{asr}=\texttt{Contriever}(\bm{\mathrm{T}}_{asr})$ (6)
DET database. While LVLMs demonstrate strong performance in object recognition, they continue to face challenges such as object counting, precise object localization, and understanding relative relationships between objects. To mitigate the issue of hallucination, which can stem from these challenges, we incorporate object detection information as auxiliary texts. We leverage a visual grounding model to extract both the object categories and their corresponding positions from sampled video frames. This approach helps provide more accurate and context-aware object detection. To enhance processing efficiency, we limit object detection to keyframes only. Specifically, we compute the CLIP similarity [29] between the object retrieval request $\bm{\mathrm{R}}_{det}$ and the sampled video frames $\bm{\mathrm{F}}$ and select relevant keyframes $\bm{\mathrm{F}}_{key}$ based on a threshold $t$:
$\bm{\mathrm{F}}_{key}=\texttt{CLIP\_similarity}(\bm{\mathrm{R}}_{det},\bm{\mathrm{F}})>t$ (7)
Once the keyframes are identified, we utilize APE [34], an efficient open-vocabulary object detection model that accepts object descriptions as prompts to detect relevant objects within frames based on specific retrieval queries. The capability of APE makes it particularly well-suited to our requirements for on-demand object retrieval in video frames. Finally, the detected objects’ categories and their corresponding positional information are stored in the DET database using natural language representations:
$DB_{det}\xleftarrow{}\bm{\mathrm{T}}_{det}=\texttt{APE}(\bm{\mathrm{F}}_{key},\bm{\mathrm{R}}_{det})$ (8)
3.4 Auxiliary Text Retrieval
During the retrieve phase, we employ the Contriever framework to encode the user’s query and the parsed requests for OCR and ASR into text embeddings, then concatenating to form the final query request $\bm{\mathrm{E}}_{req}=\texttt{Contriever}(\texttt{Concat}(\bm{\mathrm{R}},\bm{\mathrm{Q}})),\ \bm{\mathrm{R}}\in\{\bm{\mathrm{R}}_{ocr},\bm{\mathrm{R}}_{asr}\}$. Then we retrieve the auxiliary texts from $DB\in\{DB_{ocr},DB_{asr}\}$ by the FAISS tool, which computes the vector similarity between the query and text chunks stored in the database. Text chunks with a FAISS similarity score greater than threshold $t$ are indexed as the retrieval results $\bm{\mathrm{A}}\in\{\bm{\mathrm{A}}_{ocr},\bm{\mathrm{A}}_{asr}\}$. The process can be formulated as:
[…]
$\bm{\mathrm{A}}_{det}^{p}=\texttt{SceneGraph}(DB_{det})$ (10)
Finally, we acquire the object auxiliary texts based on the object information type retrieval requests $\bm{\mathrm{R}}_{type}$ of the LVLMs in the first phase, which selects and finalizes the object auxiliary information $\bm{\mathrm{A}}_{det}$. $\bm{\mathrm{A}}_{det}$ is one of the elements of the power set $\mathcal{P}$ of $\bm{\mathrm{A}}_{det}^{p}$ selected by $\bm{\mathrm{R}}_{type}$, and the retrieve process can be formulated as:
$\bm{\mathrm{A}}_{det}=\bm{\mathrm{R}}_{type}(\mathcal{P}(\bm{\mathrm{A}}_{det}^{p}))\in\mathcal{P}(\bm{\mathrm{A}}_{det}^{p})$ (11)
3.5 Integration and Generation
After obtaining different types of auxiliary texts, we organize them chronologically using natural language to create a unified auxiliary input, denoted as $\bm{\mathrm{A}}_{m}=\texttt{Concat}(\bm{\mathrm{A}}_{ocr},\bm{\mathrm{A}}_{asr},\bm{\mathrm{A}}_{det})$. These merged auxiliary inputs, along with the user’s query and the sampled video frames, are then fed into the LVLM to produce the final result. The overall process can be formulated as:
$\bm{\mathrm{O}}=\texttt{LVLM}(\bm{\mathrm{F}_{v}},\texttt{Concat}(\bm{\mathrm{A}}_{m},\bm{\mathrm{Q}}))$ (12)
4 Experiments
Model #Text LLM Params Frames Short Medium Long Overall Gain
Proprietary LVLMs
GPT-4o [28] - - 384 80.0 70.3 65.3 71.9 -
Gemini-1.5-Pro [31] - - 0.5 fps 81.7 74.3 67.4 75.0 -
Open-Source LVLMs
Video-LLaVA [16] - 7B 8 44.6 38.3 35.8 39.6 -
Video-LLaVA + Video-RAG 2.0K 7B 8 49.5 43.0 42.5 45.0 +5.4
LLaVA-NeXT-Video [48] - 7B 16 49.4 43.0 36.7 43.0 -
LLaVA-NeXT-Video + Video-RAG 2.0K 7B 16 56.6 47.4 46.0 50.0 +7.0
LongVA [47] - 7B 32 60.9 49.3 44.0 51.4 -
LongVA + Video-RAG 1.8K 7B 32 65.4 59.1 55.7 60.1 +8.7
Long-LLaVA [43] - 7B 32 60.3 51.4 44.1 52.0 -
Long-LLaVA + Video-RAG 1.9K 7B 32 66.4 60.2 59.8 62.1 +10.1
Qwen2-VL [38] - 72B 32 75.0 63.3 56.3 64.9 -
Qwen2-VL + Video-RAG 2.1K 72B 32 77.4 70.2 71.0 72.9 +8.0
LLaVA-Video [49] - 72B 32 78.0 63.7 59.6 67.1 -
LLaVA-Video + Video-RAG 2.1K 72B 32 81.1 72.9 73.1 75.7 +8.6
Table 1: Performance on the Video-MME [6] benchmark. #Text donates the average token number of auxiliary texts when inferring a single case. By applying our Video-RAG to six LVLMs, we observed an average performance improvement of 8.0% only with the addition of token counts from approximately 14 video frames (144 tokens per frame). In particular, when applying Video-RAG with 72B LLaVA-Video [49], we perform better than the proprietary method Gemini-1.5-Pro [31]. All open-source results are our replication.
4.1 Datasets
Video-MME [6] is a widely used benchmark for assessing the ability of LVLMs to handle detailed videos in real-world scenarios. It is divided into three subsets based on video length, with durations ranging from 11 seconds to 1 hour.
MLVU [52] is a long video understanding benchmark with a large wide of 9 distinct tasks. It is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours with about 12 minutes average video length.
LongVideoBench [41] is a benchmark designed to accurately retrieve and reason over detailed multimodal information from long videos, with 6,678 human-annotated multiple-choice questions in 17 fine-grained categories.
4.2 Implementation Details
We performed all experiments on NVIDIA A100 80G GPUs. During the auxiliary text generation phase, we first filter the detection requests $\bm{\mathrm{R}}_{det}$ generated by the LVLM to ensure they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. In the auxiliary text retrieval phase, we set both the CLIP similarity threshold and the FAISS similarity threshold $t$ to 0.3. We employ the IndexFlatIP as the similarity calculating method of FAISS [11]. We utilize Long-LLaVA-7B [43] for ablation studies. Since Long-LLaVA is easy to implement while supporting longer context windows, we can investigate the impact of similarity threshold selection in RAG and sampled frame rate on performance. Note that we don’t include the GPT-based Agent methods for comparison due to their resource-intensive nature (complete execution of Video-MME [6] costs around \$2000 for API purchasing when using VideoAgent [4]). Still, we include a mini-experiment of VideoAgent in the Appendix that compares the overall performance, inference time, and GPU requirements with two long-context LVLMs and Video-RAG.
4.3 Main Results
Video-MME. We evaluate our Video-RAG in four 7B open-source LVLMs, including Video-LLaVA [16], LLaVA-NeXT-Video [48], LongVA [47], Long-LLaVA [43], and two 72B LVLM Qwen2-VL [38] and LLaVA-Video [49]. To ensure a fair assessment of the performance improvements introduced by our pipeline, and given the resource constraints (especially 72B LVLMs), we reproduced the performance of all open-source models under the 32-frame setting. Results are shown in Table 1. Specifically, after applying our Video-RAG in 72B LVLM, we perform better than the SOTA proprietary model Gemini-1.5-Pro [31] (75.7% vs. 75.0%). Across the six LVLMs used in our experiments, we gained an average performance boost of 8.0%, especially a significant gain on long videos, demonstrating our pipeline’s effectiveness. This performance improvement is achieved by incorporating token counts from approximately 14 additional video frames (equivalent to 2.0K tokens), each contributing around 144 tokens under most LVLM configurations. We obtain such a large performance enhancement because most LVLMs are pre-trained primarily within the text space and aligned with visual information, often lacking explicit alignment between embedding spaces. Auxiliary texts can serve as semantic supplements sensitive to LVLMs, facilitating model activation and easing the understanding of complex videos.
MLVU. We evaluate Video-RAG when integrating into the 7B and 72B LLaVA-Video [49] of MLVU [52] in a multiple-choice task. As shown in Table 2, when applying Video-RAG in the 7B model, we achieve state-of-the-art performance at scales smaller than 72B using only a 7B model, outperforming the 32B Qryx-1.5 [24] model by 0.1%. Additionally, the 72B LLaVA-Video also has a performance gain of 0.7%. The limited performance gain may be due to having reached a bottleneck of model capacity.
Model #Params Frames Overall
Proprietary LVLMs
GPT-4o [28] - 0.5 fps 64.6
Open-Source LVLMs
Video-CCAM [5] 14B 96 63.1
Video-XL [35] 7B 256 64.9
Aria [13] 25.3B 256 70.6
LLaVA-Video* [49] 7B 64 70.8
Oryx-1.5 [24] 32B 128 72.3
LLaVA-Video* [49] 72B 64 73.1
LLaVA-Video + Video-RAG 7B 64 72.4
LLaVA-Video + Video-RAG 72B 64 73.8
Table 2: The overall performance in the multiple-choice task of the MLVU [52] benchmark. * donates the results of our replication.
LongVideoBench. We evaluate Video-RAG when applied in the 7B and 72B LLaVA-Video [49] of LongVideoBench [41]. We omit the interleaved input format introduced in LongVideoBench when applying Video-RAG. The evaluation results in Table 3 demonstrate that 72B LLaVA-Video with our Video-RAG achieves an overall performance of 65.4% on the validation set. This result surpasses the proprietary LVLM Gemini-1.5-Pro [31] by 1.4%, securing the second place, just 1.3% behind GPT-4o [28]. Meanwhile, the 7B LLaVA-Video also has a performance enhancement of 2.1% when equipped with our Video-RAG.
Model #Params Frames Overall
Proprietary LVLMs
Gemini-1.5-Pro [31] - 256 64.0
GPT-4o [28] - 256 66.7
Open-Source LVLMs
VideoChat2-Mistral [14] 7B 8 39.3
ShareGPT4Video [2] 7B 8 39.7
LLaVA-Next-Mistral [20] 7B 8 49.1
PLLaVA [42] 34B 16 53.2
LLaVA-Video* [49] 7B 64 56.6
LLaVA-Video* [49] 72B 64 61.9
LLaVA-Video + Video-RAG 7B 64 58.7
LLaVA-Video + Video-RAG 72B 64 65.4
Table 3: The overall performance on the validation set of LongVideoBench [41]. * donates the results of our replication.
4.4 Ablation Studies
Effect of different sampling frame number. To explore the effect of the number of sampling frames on Video-RAG, we experience sampling frames number 8, 16, 32, and 64 frames in 7B model Long-LLaVA [43], results are shown in Figure 4. As demonstrated, Video-RAG consistently delivers performance improvements across all frame rates especially in long videos, with these gains increasing as the frame rate rises. Furthermore, the results indicate that the highest accuracy without Video-RAG is achieved when 32 frames are sampled from the video. Therefore, we adopt this configuration in the subsequent ablation experiments.
Figure 4: Performance with different sampling frames rate on Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM.
Effect of different components of Video-RAG. To explore the effectiveness of auxiliary texts, we incrementally add object detection, OCR, and ASR as auxiliary texts after retrieving by the RAG system to evaluate Long-LLaVA-7B [43] in the Video-MME [6] benchmark. As shown in Table 4, the performance of Long-LLaVA progressively improves as auxiliary texts are incrementally added (52.0% → 52.9% → 54.3% → 62.1%). Among these components, ASR auxiliary texts contribute to a general improvement for different video durations, underscoring the critical role of audio transcription that can provide comprehensive information beyond visual cues in video understanding. Meanwhile, we randomly sampled an equivalent token number of auxiliary texts to serve as inputs for assessing the effectiveness of RAG retrieval. The experiment demonstrated a 3.0% improvement in performance after incorporating RAG. We also evaluate the performance across sub-tasks within Video-MME [6] and other video benchmarks like VNBench [51], more details are shown in the Appendix.
Figure 5: Qualitative result shown in Video-MME [6] benchmark when applying Video-RAG with LLaVA-Video [49].
Figure 6: Grad-CAM visualizations of the last hidden state heatmap along with t-SNE visualizations of the user’s query and keyframe features of the example shown in Figure 5. As demonstrated, the retrieved auxiliary texts help cross-modality alignment by assisting the model to pay more attention to query-relevant keyframes and thus generate more robust and accurate answers to the user’s query.
RAG DET OCR ASR Short Medium Long Overall
60.3 51.4 44.1 52.0
66.2 54.7 50.3 57.1
61.4 51.9 45.2 52.9
63.2 53.2 46.3 54.3
65.7 55.8 56.0 59.1
66.4 60.2 59.8 62.1
Table 4: Results on combinations of different auxiliary texts in Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM.
Effect of different threshold of RAG processing. When using the RAG tool for retrieval, we specify a similarity threshold $t$ as a criterion for information selection. In the retrieval for OCR and ASR texts, information is selected if its FAISS similarity exceeds $t$. For object detection, frames are selected as keyframes based on their CLIP similarity surpassing $t$, and the relevant information is then extracted. Setting $t$ too high may hinder the retrieval of relevant information while setting it too low can result in information redundancy and increased reasoning complexity. To investigate this trade-off, we conduct ablation experiments to evaluate the impact of different threshold values. The results are shown in Table 5. Notably, $t=0$ and $t=1$ correspond to all auxiliary texts input into the model and no auxiliary texts input, respectively. To balance performance with information density and processing time (especially APE [34] detection in keyframes), we selected a threshold of 0.3 for our implementation. More details about similarity scores are shown in the Appendix. Under this configuration, the additional text length of approximately 1.9K tokens typically remains within the context window limits of open-source LVLMs. For models with more stringent context window limitations, a threshold of 0.4 may also be a viable option.
$t$Token Time Short Medium Long Overall
0.0 3.6K 36s 67.6 59.4 59.1 62.0
0.1 3.4K 30s 67.0 59.7 59.1 61.9
0.2 2.7K 18s 66.0 60.2 59.2 61.8
0.3 1.9K 11s 66.4 60.2 59.8 62.1
0.4 0.8K 8s 65.6 58.0 58.3 60.6
0.5 0.3K 7s 63.1 54.9 50.2 56.1
1.0 0.0K 6s 60.3 51.4 44.1 52.0
Table 5: Performance with different thresholds of retrieval on Video-MME [6] when using Long-LLaVA-7B [43] as the LVLM. #Token and Time denote the total token number of the auxiliary texts and the average inference time per question, respectively.
4.5 Qualitative Evaluation
We present qualitative results in the case of Video-MME [6] in Figure 5 and Figure 6. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate responses to user queries. Grad-CAM [32] and t-SNE [37] visualization results also show that applying Video-RAG helps the LVLM’s cross-modality alignment.
5 Conclusion
In this paper, we present Video-RAG for effective long video understanding by integrating retrieved auxiliary texts with LVLMs, achieving proprietary-level performance with 72B open-source LVLM. Unlike traditional long-context and GPT-based Agent methods that may have limited performance gain and are resource-intensive, Video-RAG offers a resource-efficient, plug-and-play solution for any open-source LVLMs that only leverage open-source external tools to extract the visually-aligned auxiliary texts from pure video data for input. In the future, we will explore how to more efficiently integrate auxiliary texts and provide an adaptive frame selection strategy for LVLMs.
Supplementary Material
6 Decouple Query
In the initial phase of the proposed Video-RAG, we employ a decouple prompt, denoted as $\bm{\mathrm{P}}$, to guide the LVLM in generating retrieval requests. In this section, we present one example of a prompt designed for multiple-choice questions, as illustrated in Figure 9.
7 Sub-set of Video-MME
As outlined in the implementation details, we randomly sampled a subset of the Video-MME [6] dataset to evaluate a computationally resource-intensive, agent-based method with long-context LVLMs. Specifically, we selected 10% of the full dataset, comprising 30 short, 30 medium-length, and 30 long videos. Each video contains three multiple-choice questions. Importantly, we ensured that the performance ranking of the methods on the subset mirrored that of the full dataset. As shown in Tables 6 and 7, we evaluated four distinct 7B models Chat-Univi-v1.5 [10], LLaVA-NeXT-Video [48], LongVA [47], and Long-LLaVA [43] using a frame sampling rate of 16 for both the subset and the full set. Our results indicate that the performance rankings remained consistent across both evaluations.
Method Short Medium Long Overall
Chat-Univi-v1.5 [10] 50.0 33.3 17.8 33.7
LLaVA-NeXT-Video [48] 54.4 33.3 23.3 37.0
LongVA [47] 56.7 50.0 38.9 48.5
Long-LLaVA [43] 58.9 52.2 40.0 50.4
Table 6: Performance of Video-MME sub-set.
Method Short Medium Long Overall
Chat-Univi-v1.5 [10] 45.7 39.0 35.7 40.1
LLaVA-NeXT-Video [48] 51.1 41.8 36.8 43.2
LongVA [47] 60.8 45.2 41.4 49.1
Long-LLaVA [43] 59.3 49.3 44.4 51.0
Table 7: Performance of Video-MME full-set.
8 Results on Video-MME Sub-Set
We examine Video-RAG against two representative methods in terms of inference time, GPU resource requirements, and overall performance. Given that GPT-based Agent methods are resource-intensive, we randomly sampled a sub-set of the Video-MME [6] for evaluation, as described in Section 7. As demonstrated in Figure 7, VideoAgent [4], a typically GPT-based Agent method, requires significant time to process video and deliver suboptimal performance. Meanwhile, LongVA [47], a representative long-context LVLM, shows limited improvement from increasing the frame rate and even experiences performance degradation. Integrating our Video-RAG into the 16-frame LongVA results in substantial performance improvements while reducing GPU resource consumption. Specifically, with only increasing 8GB GPU memory compared to the base (16-frames LongVA), we achieve 11.5% overall performance improvement, while outperforming another long-context LVLM Long-LLaVA-7B [43] in 16-frames setting by 9.6% with less GPU memory requirements and compatible total inference time. These results demonstrated that our Video-RAG is lightweight with lower computing overhead than the other typical methods.
Figure 7: The comparison of our Video-RAG with two common approaches. The size of the bubbles represents the total time consumed for completing inference on the Video-MME [6] sub-set.
9 Details of Similarity Score Calculation
In the process of using the RAG system to retrieve auxiliary texts extracted from videos, we define a similarity threshold $t$ to ensure the selection of relevant texts. Specifically, we employ FAISS-based [11] similarity to select OCR and ASR texts, while CLIP [29] similarity is used for keyframe selection. In our implementation, the similarity threshold $t$ is set to 0.3. As for OCR and ASR selection, For any given list of the retrieve request $\bm{\mathrm{R}}$ and auxiliary texts $\bm{\mathrm{A}}$, the Contriever [8] framework maps the text to a text embedding as:
$\bm{\mathrm{E}}_{a_{i}}=\texttt{Contriever}(\bm{\mathrm{A}}_{i}),\quad i=1,2,\dots,n$
$\bm{\mathrm{E}}_{r_{i}}=\texttt{Contriever}(\bm{\mathrm{R}}_{i}),\quad i=1,2,\dots,n$
The average embedding of the retrieve request is then computed as:
$\bm{\mathrm{E}}_{r}=\frac{1}{n}\sum_{i=1}^{n}\bm{\mathrm{E}}_{r_{i}}$
After that, the embedding of the request and the list of auxiliary texts is normalized:
$\bm{\mathrm{E}}_{a_{i}}=\frac{\bm{\mathrm{E}}_{a_{i}}}{\|\ \bm{\mathrm{E}}_{a_{i}}\|},\quad\bm{\mathrm{E}}_{r}=\frac{\bm{\mathrm{E}}_{r}}{\|\ \bm{\mathrm{E}}_{r}\|}$
The similarity between the query embedding $\bm{\mathrm{E}}_{r}$ and the document vector $\bm{\mathrm{E}}_{a}$ is computed using the inner product, the FAISS library is employed to efficiently perform this search and return the indices of the auxiliary texts meeting the criterion:
$S(\bm{\mathrm{E}}_{r},\bm{\mathrm{E}}_{a_{i}})=\bm{\mathrm{E}}_{r}\cdot\bm{\mathrm{E}}_{a_{i}}>t$
As for object detection, we use CLIP to select the video keyframe. During this process, we first filter the object detection request $\bm{\mathrm{R}}_{det}$ to ensure they correspond to CLIP-sensitive physical entities, avoiding the inclusion of abstract concepts. Specifically, if it is a single word, direct part-of-speech filtering is applied; if it is a compound word, certain rules are followed to check for compliance, such as whether it is an adjective plus a noun, or a noun plus a noun. We use the Spacy library to achieve this. After this, we put the text “A picture of” before each object detection request.
Then, we extracting embedding from both the video frames $\bm{\mathrm{F}}$ and the detection request $\bm{\mathrm{R}}_{det}$:
$\bm{\mathrm{E}}_{\bm{\mathrm{F}}_{j}}=\texttt{CLIP}(\bm{\mathrm{F}}_{j}),\quad j=1,2,\dots,m$
$\bm{\mathrm{E}}_{\bm{\mathrm{R}}_{i}}=\texttt{CLIP}(\bm{\mathrm{R}}_{det_{i}}),\quad i=1,2,\dots,n$
The similarity between each video frame and the detection retrieve requests is computed using the dot product between the image and text feature embeddings. For each frame $\bm{\mathrm{F}}_{j}$, and for each retrieve request $\bm{\mathrm{E}}_{\bm{\mathrm{R}}_{i}}$, the similarity score is given by:
$S_{ij}=\bm{\mathrm{E}}_{\bm{\mathrm{F}}_{j}}\cdot\bm{\mathrm{E}}_{\bm{\mathrm{R}}_{i}}$
where $\cdot$ denotes the dot product. The final similarity score for each frame is the average similarity across all requests:
$S_{j}=\frac{1}{n}\sum_{i=1}^{n}S_{ij}$
This computes the mean similarity for each frame across all text descriptions, resulting in a similarity vector $\mathbf{S}=[S_{1},S_{2},\dots,S_{m}]$. The similarity scores are adjusted by a scaling factor $\alpha$, which is computed based on the number of frames $m$ and a base frame number $b$ (which is set to 16 and 4.0, respectively) to adapted different video sampling rate of LVLMs:
$\alpha=\beta\times\frac{m}{b}$
where $\beta$ is a predefined scaling parameter.
Next, the similarity scores are scaled and normalized to ensure that they sum to 1:
$S_{j}^{\text{norm}}=\frac{\alpha\times S_{j}}{\sum_{k=1}^{m}S_{k}}$
where $S_{j}^{\text{norm}}$ represents the normalized similarity score for frame $\bm{\mathrm{F}}_{j}$.
The final step is to select the keyframes based on the normalized similarity scores. A threshold $t$ is applied to the normalized similarities, such that frames with similarity scores above the threshold are selected as keyframes:
[…]
Thus, the set of selected keyframes is given by:
$\bm{\mathrm{F}}_{key}=\{\bm{\mathrm{F}}_{j}\mid S_{j}^{\text{norm}}>t,\,j=1,2,\dots,m\}$
10 More Ablation Studies
Effect of different components of Video-RAG. We evaluate the performance across sub-tasks within Video-MME [6], as shown in Figure 8. The results reveal that object detection auxiliary texts significantly enhance spatial perception and object counting, while OCR auxiliary texts specifically improve performance on text recognition tasks. Additionally, ASR auxiliary texts contribute to a general improvement in inference tasks, underscoring the critical role of audio transcription in video understanding. Given that audio transcription is considerably more time-consuming than character recognition or object detection, these texts should be selected based on the requirements of the application.
Figure 8: Performance on 12 sub-tasks in Video-MME [6] benchmark after applying different components in Long-LLaVA.
Besides studying the inference of different components of Video-RAG in the Video-MME [6] benchmark, we also experiment with a different type of video benchmark VNBench [51] with Long-LLaVA-7B [43]. VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. VNBench randomly inserts stickers or text into the video that has nothing to do with the original content of the video, thus typically challenging the model’s needle-in-the-haystack capability. As shown in Table 8, we find that applying DET and OCR as auxiliary texts can significantly improve the performance in retrieval, ordering, and counting tasks. However, the ASR component will decline the performance due to the subtitles are not ancillary to this particular task. These results demonstrated that our proposed distinct types of auxiliary texts can be selected according to the application needs to meet the requirements better.
RAG DET OCR ASR Ret Ord Cnt Overall
65.1 25.6 24.2 38.3
66.9 28.4 23.8 39.7
68.2 31.3 28.9 42.8
66.7 31.3 29.6 42.5
Table 8: Results on combinations of different auxiliary texts in VNBench [51] with 1-try setting when applying 7B Long-LLaVA [43] as LVLM under the 32-frames setting. Ret, Ord, and Cnt represent retrieval, ordering, and counting tasks, respectively.
11 More Qualitative Results
In this section, we show more results of LLaVA-Vdieo-7B when applying Video-RAG in different examples in Figure 10. The figure highlights several representative cases involving detailed video comprehension from Video-MME [6]. As illustrated, augmenting LLaVA-Video with external tools to process and retrieve auxiliary texts from videos significantly enhances its ability to reduce visual hallucinations, thereby enabling more accurate and confident responses to user queries.
Figure 9: Decouple prompt of the multiple-choice question for LVLMs.
Figure 10: Qualitative results of LLaVA-Vdieo when applying Video-RAG.